# Learning Where to Learn: Training Distribution Selection for Provable OOD Performance

### v1

@@ 4 Practical algorithms for training distribution design
$\inf_{\nu\in\mathscr{P}_{2}(\mathcal{U})}{\mathopen{}\mathclose{{}\left\{\mathscr{E}_{\mathbb{Q}}\bigl{(}\widehat{\mathcal{G}}^{(\nu)}\bigr{)}\,\middle|\,\widehat{\mathcal{G}}^{(\nu)}\in\operatorname*{arg\,min}_{\mathcal{G}\in\mathcal{H}}\operatorname{\mathbb{E}}_{u\sim\nu}\big{\lVert}\mathcal{G}^{\star}(u)-\mathcal{G}(u)\big{\rVert}_{\mathcal{Y}}^{2}}\right\}}\,,$ (4.1)
where $\mathscr{E}_{\mathbb{Q}}$ is the average-case OOD accuracy functional (3.3) and $\mathcal{H}$ is the model hypothesis class. Eqn. (4.1) is a bilevel optimization problem. The inner minimization problem seeks the best-fit model $\widehat{\mathcal{G}}^{(\nu)}$ for a fixed training distribution $\nu$. The outer problem minimizes $\mathscr{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\nu)})$ over $\nu$ in the space of probability measures. The map $\nu\mapsto\widehat{\mathcal{G}}^{(\nu)}$ can be highly nonlinear, which poses a challenge to solve (4.1). To devise practical solution methods, this section considers the bilevel optimization problem (4.1) in Sec. 4.1 and a computationally tractable relaxation of (4.1) in Sec. 4.2.
@@ 4.1 Exact bilevel formulation
This subsection focuses on directly solving the bilevel optimization problem (4.1) to identify the optimal training distribution. Bilevel optimization has been extensively studied in the literature; see, for example, [13, 24, 53]. For the infinite-dimensional problem (4.1), our approach begins by specifying a Hilbert space $\mathcal{H}$. Using its linear structure, we calculate the gradient of the Lagrangian associated with (4.1) with respect to the variable $\nu$. This gradient computation leverages the adjoint state method and the envelope theorem [3]. In what follows, we specialize this general framework to the case where $\mathcal{H}$ is a reproducing kernel Hilbert space (RKHS) with scalar output space $\mathcal{Y}=\mathbb{R}$. The resulting algorithm can be efficiently implemented using representer theorems for kernel methods [38]. Additionally, other linear classes of functions can be incorporated into this framework with minimal effort. Extending the framework to nonlinear, vector-valued function spaces $\mathcal{H}$, such as neural network spaces, presents an exciting direction for future research.
@@
Assume that $\mathcal{G}^{\star}\in L^{2}_{\nu}\equiv L^{2}_{\nu}(\mathcal{U};\mathbb{R})$ for every $\nu\in\mathscr{P}_{2}(\mathcal{U})$ under consideration; this is true, for example, if $\mathcal{G}^{\star}$ is uniformly bounded on $\mathcal{U}$. Define the data misfit $\Psi\colon\mathcal{H}\to\mathbb{R}$ by $\Psi(\mathcal{G})\coloneqq\frac{1}{2}\lVert\iota_{\nu}\mathcal{G}-\mathcal{G}^{\star}\rVert^{2}_{L^{2}_{\nu}}$; this is an abstract way to write the convex inner objective in (4.1) (rescaled by $1/2$). By Lem. C.1, the inner solution $\widehat{\mathcal{G}}^{(\nu)}$ solves the equation $\mathcal{K}_{\nu}\widehat{\mathcal{G}}^{(\nu)}=\iota^{*}_{\nu}\mathcal{G}^{\star}$ in $\mathcal{H}$. This fact allows us to define the Lagrangian $\mathcal{L}\colon\mathscr{P}_{2}(H)\times\mathcal{H}\times\mathcal{H}\to\mathbb{R}$ for (4.1) by
@@
SM C.1 provides the proofs of Lem. 4.1 and Prop. 4.2.
@@
A nonparametric approach to solve the bilevel problem (4.1) is to derive the Wasserstein gradient flow and its Lagrangian equivalence $\dot{u}=-\nabla(D\mathcal{J})(\operatorname{Law}(u))(u)$, where $D\mathcal{J}$ is as in Prop. 4.2. This leads to a curve of measures $t\mapsto\operatorname{Law}(u(t))$ converging to a steady-state probability distribution. But this dynamic requires gradients of the true map $\mathcal{G}^{\star}$, which may not be available or could be too expensive to compute even if available. Although beyond the scope of this paper, future work will discretize this mean-field evolution with interacting particles.
@@
Although Prop. 4.2 provides the desired derivative of the bilevel problem (4.1), it does not specify how to use the derivative to compute the optimal $\nu$. This issue is nontrivial because $\nu\in\mathscr{P}_{2}(\mathcal{U})$ belongs to a nonlinear metric space. One approach assumes a parametrization $\nu=\nu_{\vartheta}$, where $\vartheta\in\mathcal{V}$ for some separable Hilbert parameter space $\mathcal{V}$. Typically $\mathcal{V}\subseteq\mathbb{R}^{P}$ for some $P\in\mathbb{N}$. We then use the chain rule to apply Euclidean gradient-based optimization tools to solve (4.1). To this end, define $\mathsf{J}\colon\mathcal{V}\to\mathbb{R}$ by $\vartheta\mapsto\mathsf{J}(\vartheta)\coloneqq\mathcal{J}(\nu_{\vartheta})$, where $\mathcal{J}$ is as in Prop. 4.2. The following theorem computes the gradient $\nabla\mathsf{J}\colon\mathcal{V}\to\mathcal{V}$.
@@
In Thm. 4.4, $\mu_{0}$ is typically the Lebesgue measure if $\mathcal{U}\subseteq\mathbb{R}^{d}$. SM C.1 contains the proof.
@@
The parameters $\vartheta$ of $\nu_{\vartheta}$ can now be updated using any gradient-based optimization method. For instance, a simple Hilbert gradient flow is given by $\dot{\vartheta}=-\nabla\mathsf{J}(\vartheta)$.
@@
A gradient descent scheme for the bilevel optimization problem (4.1) is summarized in Alg. 1. Alternative gradient-based optimization methods can also be explored. Discretization details for the gradient (4.4) are provided in SM C.1.
@@ 4.2 Upper-bound minimization
It is thus natural to approximate the solution to (4.1) by the minimizing distribution of the right-hand side of (4.6). The dependence of the factor $c$ in Prop. 3.1 on the Lipschitz constant of $\widehat{\mathcal{G}}^{(\nu)}$ is hard to capture in practice. One way to circumvent this difficulty is to further assume an upper bound $\operatorname{Lip}(\widehat{\mathcal{G}}^{(\nu)})\leq R$ uniformly in $\nu$ for some constant $R>0$. We define $c_{R}(\mathcal{G}^{\star},\nu,\nu^{\prime})$ to be $c(\mathcal{G}^{\star},\widehat{\mathcal{G}}^{(\nu)},\nu,\nu^{\prime})$ with all instances of $\operatorname{Lip}(\widehat{\mathcal{G}}^{(\nu)})$ replaced by $R$. The assumption of a uniform bound on the Lipschitz constant is valid for certain chosen model classes $\mathcal{H}$, e.g., kernels with Lipschitz feature maps or Lipschitz-constrained neural networks [25].
@@
The proof is straightforward. Line 3 ensures that the first term of the objective (4.7) does not increase, while keeping the second term constant. Next, Line 4 minimizes both terms. Iterating this argument, we deduce that the algorithm maintains a non-increasing objective value for all $k$. ∎
@@
If the set $\mathsf{P}$ contains distributions with uniformly bounded second moments and the first term on the right-hand side of (4.6) is uniformly bounded over $\mathsf{P}$, then the optimal value of (4.7) is bounded above by the value of another minimization problem whose solution is the $2$-Wasserstein barycenter $\nu_{\mathbb{Q}}\in\mathscr{P}_{2}(\mathcal{U})$ of $\mathbb{Q}\in\mathscr{P}_{2}(\mathscr{P}_{2}(\mathcal{U}))$ [7, Thm 3.7, p. 444]. An analogous bound is even easier to verify for the $\mathsf{W}_{1}$ result (3.1). Put another way, the $\mathcal{G}^{\star}$-dependent solution of the proposed problem (4.7) over $\mathsf{P}$ (and hence also the solution of Eqn. (4.1)) is always at least as good as the barycenter $\nu_{\mathbb{Q}}$, which only depends on $\mathsf{P}$ and $\mathbb{Q}$.
@@
and $f\colon\mathcal{U}\to\mathbb{R}$. The next result differentiates $F$; the proof is in SM C.2.
@@
Taking the gradient of $u\mapsto DF(\nu)(u)$ in (4.9) yields the Wasserstein gradient, which can be utilized in Wasserstein gradient flow discretizations; see Rmk. 4.3. In the case of parametric distributions, combining Prop. 4.8 with the chain rule yields gradient descent schemes similar to those introduced in Sec. 4.1. For further details, see Lem. C.8 and its special case derived from Rmk. C.9 in SM C.2.
@@ 5 Numerical results
This section implements Alg. 1 (the bilevel gradient descent algorithm) for function approximation using kernel methods and Alg. 2 (the alternating minimization algorithm) for two operator learning tasks. Comprehensive experiment details and settings are collected in SM E.
@@
We instantiate a discretization of Alg. 1 on several function approximation benchmarks. The test functions $\mathcal{G}^{\star}\in\{g_{i}\}_{i=1}^{4}$ each map $\mathbb{R}^{d}$ to $\mathbb{R}$ and are comprised of the Ishigami ($g_{1}$), Sobol G ($g_{2}$), and Friedmann functions ($g_{3},g_{4}$). We use the kernel framework from Sec. 4.1 with $\mathcal{H}$ the RKHS of the squared exponential kernel with a fixed scalar bandwidth hyperparameter. Also, $\mathbb{Q}$ is an empirical measure with $10$ atoms centered at fixed Gaussian measures realized by standard normally distributed means and Wishart distributed covariance matrices. A cosine annealing scheduler determines the gradient descent step sizes. The training distribution $\nu_{\vartheta}\coloneqq\mathcal{N}(m_{0},\mathcal{C})$ is parametrized by the covariance matrix $\mathcal{C}$, which we optimize using Alg. 1.
@@
Table 1: Function approximation with bilevel Alg. 1. Reported values are mean $\mathsf{Err}$ for kernel regressors trained with $2^{14}$ labeled data samples from (i) the initial training distribution $\mathcal{N}(m_{0},I_{d})$ and (ii) the optimal $\mathcal{N}(m_{0},\widehat{\mathcal{C}})$ from Alg. 1. One standard deviation from the mean $\mathsf{Err}$ over $10$ independent runs is also reported.
$g_{1}~{}(d=3)$ $g_{2}~{}(d=2)$ $g_{3}~{}(d=5)$ $g_{3}~{}(d=8)$ $g_{4}~{}(d=4)$ $g_{4}~{}(d=5)$
$\mathcal{N}(m_{0},I_{d})$ $0.980\pm 0.001$ $0.688\pm 0.009$ $0.754\pm 0.014$ $0.919\pm 0.004$ $0.892\pm 0.006$ $0.902\pm 0.004$
$\mathcal{N}(m_{0},\widehat{\mathcal{C}})$ $0.872\pm 0.006$ $0.016\pm 0.000$ $0.144\pm 0.002$ $0.442\pm 0.003$ $0.104\pm 0.009$ $0.320\pm 0.007$
In Table 1, we report the root relative average OOD squared error $\mathsf{Err}\coloneqq(\mathscr{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\nu)})/\mathscr{E}_{\mathbb{Q}}(0))^{1/2}$ for models trained on the initial and final iterates $\nu_{\vartheta}$ of Alg. 1. However, the performance gap can close depending on the complexity of $\mathbb{Q}$. Fig. 2 specializes to the $g_{3}$ case. Although fitting the model on data sampled from the optimal $\nu_{\vartheta}$ seems to uniformly outperform standard normal distribution sampling, when corrected for the “offline” cost of finding the optimal $\nu_{\vartheta}$, the gains are more modest.
@@
Figure 2: Alg. 1 applied to ground truth $g_{3}\colon\mathbb{R}^{5}\to\mathbb{R}$. (Left) Evolution of $\mathsf{Err}$ over $1000$ iterations of gradient descent. (Center) $\mathsf{Err}$ of model trained on $N$ samples from initial $\nu_{\vartheta}$ vs. optimal $\nu_{\vartheta}$. (Right) Same as center, except incorporating the additional function evaluation cost incurred from Alg. 1. Shading represents two standard deviations away from the mean over $10$ independent runs.
@@
We apply the alternating scheme Alg. 2 to two important operator learning problems in the computational sciences. Let $\Omega\subset\mathbb{R}^{2}$ denote a bounded domain with boundary $\partial\Omega$. Consider the elliptic PDE $\nabla\cdot(a(x)\nabla u(x))=0$, where $x\in\Omega$, $a=a(x)>0$ is the conductivity, and $u=u(x)$ is the PDE solution. Given a mean-zero Neumann boundary condition (BC) $g=a\frac{\partial u}{\partial\textbf{n}}|_{\partial\Omega}$, one can obtain a unique PDE solution $u$ and the resulting Dirichlet data $f\coloneqq u|_{\partial\Omega}$. For a fixed conductivity $a$, we define a Neumann-to-Dirichlet (NtD) map $\Lambda_{a}^{-1}\colon g\mapsto f$, which is also used in the EIT problem [18]; see also Sec. 1. On the other hand, the map $\mathcal{G}\colon a\mapsto u$ with a fixed Neumann or Dirichlet BC (and nonzero source) is the Darcy flow parameter-to-solution operator, which plays a crucial role in geothermal energy extraction and carbon storage applications. Learning surrogate operators for $\Lambda_{a}^{-1}$ and $\mathcal{G}$ can enable rapid simulations for efficient inference and control.
@@
Figure 3: (Left) Conductivity for NtD learning. (Middle) After $80$ independent runs, decay of the average relative OOD error of the model when trained on the optimal distribution identified at each iteration of Alg. 2; a $95$% confidence interval of the true relative OOD error is provided at each iteration. (Right) Decay of average AMA loss defined in (4.7) vs. iteration relative to the same loss at initialization.

### v3

@@ 4 Practical algorithms for training distribution design
$\inf{\mathopen{}\mathclose{{\left\{\mathscr{E}_{\mathbb{Q}}\bigl(\widehat{\mathcal{G}}^{(\nu)}\bigr)\,\middle|\,\nu\in\mathscr{P}_{2}(\mathcal{U})\quad\text{and}\quad\widehat{\mathcal{G}}^{(\nu)}\in\operatorname*{arg\,min}_{\mathcal{G}\in\mathcal{H}}\operatorname{\mathbb{E}}_{u\sim\nu}\big\lVert\mathcal{G}^{\star}(u)-\mathcal{G}(u)\big\rVert_{\mathcal{Y}}^{2}}}\right\}}\,$, (4.1)
where $\mathscr{E}_{\mathbb{Q}}$ is the average-case OOD accuracy functional (3.3) and $\mathcal{H}$ is the model hypothesis class.
Eqn. (4.1) is a bilevel optimization problem. The inner minimization problem seeks the best-fit model $\widehat{\mathcal{G}}^{(\nu)}$ for a fixed training distribution $\nu$. The outer problem minimizes $\mathscr{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\nu)})$ over $\nu$ in the space of probability measures. The map $\nu\mapsto\widehat{\mathcal{G}}^{(\nu)}$ can be highly nonlinear, which poses a challenge to solve (4.1). To devise practical solution methods, this section considers the bilevel optimization problem (4.1) in Sec. 4.1 and a computationally tractable relaxation of (4.1) in Sec. 4.2.
Remark 4.1 (finite data).
It is necessary to use finite data to estimate the optimal training distribution (SM D.1). We thus assume sample access to the chosen feasible set of training distributions, e.g., i.i.d. sampling or interacting particles. Moreover, we only require sample access to samples from $\mathbb{Q}$.
@@ 4.1 Exact bilevel formulation
This subsection focuses on directly solving the bilevel optimization problem (4.1) to identify the optimal training distribution. Bilevel optimization has been extensively studied in the literature; see, for example, the work of Colson et al. [22], Golub and Pereyra [40], Van Leeuwen and Aravkin [93]. For the infinite-dimensional problem (4.1), our approach begins by specifying a Hilbert space $\mathcal{H}$. Using its linear structure, we calculate the gradient of the Lagrangian associated with (4.1) with respect to the variable $\nu$. This gradient computation leverages the adjoint state method and the envelope theorem [4]. In what follows, we specialize this general framework to the case where $\mathcal{H}$ is a reproducing kernel Hilbert space (RKHS) with scalar output space $\mathcal{Y}=\mathbb{R}$. The resulting algorithm can be efficiently implemented using representer theorems for kernel methods [65]. Additionally, other linear classes of functions can be incorporated into this framework with minimal effort. Extending the framework to nonlinear, vector-valued function spaces $\mathcal{H}$, such as neural network or neural operator spaces, is an exciting direction for future research.
@@
Assume that $\mathcal{G}^{\star}\in L^{2}_{\nu}\equiv L^{2}_{\nu}(\mathcal{U};\mathbb{R})$ for every $\nu\in\mathscr{P}_{2}(\mathcal{U})$ under consideration; this is true, for example, if $\mathcal{G}^{\star}$ is uniformly bounded on $\mathcal{U}$. Define the data misfit $\Psi\colon\mathcal{H}\to\mathbb{R}$ by $\Psi(\mathcal{G})\coloneqq\frac{1}{2}\lVert\iota_{\nu}\mathcal{G}-\mathcal{G}^{\star}\rVert^{2}_{L^{2}_{\nu}}$; this is an abstract way to write the convex inner objective in (4.1) (rescaled by $1/2$). By Lem. D.1, the inner solution $\widehat{\mathcal{G}}^{(\nu)}$ solves the equation $\mathcal{K}_{\nu}\widehat{\mathcal{G}}^{(\nu)}=\iota^{*}_{\nu}\mathcal{G}^{\star}$ in $\mathcal{H}$. This fact allows us to define the Lagrangian $\mathcal{L}\colon\mathscr{P}_{2}(H)\times\mathcal{H}\times\mathcal{H}\to\mathbb{R}$ for (4.1) by
@@
SM D.2 provides the proofs of Lem. 4.2 and Prop. 4.3 and implementation details.
@@
A nonparametric approach to solve the bilevel problem (4.1) is to derive the Wasserstein gradient flow and its Lagrangian equivalence $\dot{u}=-\nabla(D\mathcal{J})(\operatorname{Law}(u))(u)$, where $D\mathcal{J}$ is as in Prop. 4.3. This leads to a curve of measures $t\mapsto\operatorname{Law}(u(t))$ converging to a steady-state probability distribution. It is important to note that this dynamic requires gradients of the true map $\mathcal{G}^{\star}$, which may not be available or could be too expensive to compute even if available. To implement this mean-field evolution, we can discretize it with interacting particles; see SM F.5.
@@
Although Prop. 4.3 provides the desired derivative of the bilevel problem (4.1), it does not specify how to use the derivative to compute the optimal $\nu$. This issue is nontrivial because $\nu\in\mathscr{P}_{2}(\mathcal{U})$ belongs to a nonlinear metric space. One approach assumes a parametrization $\nu=\nu_{\vartheta}$, where $\vartheta\in\mathcal{V}$ for some separable Hilbert parameter space $\mathcal{V}$. Typically $\mathcal{V}\subseteq\mathbb{R}^{P}$ for some $P\in\mathbb{N}$. We then use the chain rule to apply Euclidean gradient-based optimization tools to solve (4.1). To this end, define $\mathsf{J}\colon\mathcal{V}\to\mathbb{R}$ by $\vartheta\mapsto\mathsf{J}(\vartheta)\coloneqq\mathcal{J}(\nu_{\vartheta})$, where $\mathcal{J}$ is as in Prop. 4.3. The following theorem computes the gradient $\nabla\mathsf{J}\colon\mathcal{V}\to\mathcal{V}$.
@@
In Thm. 4.5, $\mu_{0}$ is typically the Lebesgue measure if $\mathcal{U}\subseteq\mathbb{R}^{d}$. SM D.2 contains the proof.
@@
The parameters $\vartheta$ of $\nu_{\vartheta}$ can now be updated using any gradient-based optimization method. For instance, a simple gradient flow in Hilbert space is given by $\dot{\vartheta}=-\nabla\mathsf{J}(\vartheta)$.
@@
A gradient descent scheme for the bilevel optimization problem (4.1) is summarized in Alg. 1. Alternative gradient-based optimization methods can also be explored. We emphasize that the derived bilevel gradients are exact; in implementation, the only source of error arises from discretization. Discretization details for the gradient (4.4) are provided in SM D.2; in particular, we “empiricalize” all expectations.
@@ 4.2 Upper-bound minimization
It is thus natural to approximate the solution to (4.1) by the minimizing distribution of the right-hand side of (4.6). The dependence of the factor $c$ in Prop. 3.1 on the Lipschitz constant of $\widehat{\mathcal{G}}^{(\nu)}$ is hard to capture in practice. One way to circumvent this difficulty is to further assume an upper bound $\operatorname{Lip}(\widehat{\mathcal{G}}^{(\nu)})\leq R$ uniformly in $\nu$ for some constant $R>0$. We define $c_{R}(\mathcal{G}^{\star},\nu,\nu^{\prime})$ to be $c(\mathcal{G}^{\star},\widehat{\mathcal{G}}^{(\nu)},\nu,\nu^{\prime})$ with all instances of $\operatorname{Lip}(\widehat{\mathcal{G}}^{(\nu)})$ replaced by $R$. The assumption of a uniform bound on the Lipschitz constant is valid for certain chosen model classes $\mathcal{H}$, e.g., kernels with Lipschitz feature maps or Lipschitz-constrained neural networks [41]. Although such Lipschitz control in deep neural networks is challenging, there has been substantial recent progress on architectures with explicit Lipschitz constraints, e.g., spectral normalization, orthogonal/Householder layers, 1-Lipschitz activations [66, 7, 95, 70]. Thus, expressive neural models with controlled Lipschitz behavior can be made practical.
@@
Line 3 ensures that the first term of the objective (4.7) does not increase, while keeping the second term constant. Next, Line 4 minimizes both terms. Iterating this argument, we deduce that the algorithm maintains a non-increasing objective value for all $k$. ∎
@@
If the set $\mathsf{P}$ contains distributions with uniformly bounded second moments and the first term on the right-hand side of (4.6) is uniformly bounded over $\mathsf{P}$, then the optimal value of (4.7) is bounded above by the value of another minimization problem whose solution in $\mathscr{P}_{2}(\mathcal{U})$ is the $2$-Wasserstein barycenter of $\mathbb{Q}\in\mathscr{P}_{2}(\mathscr{P}_{2}(\mathcal{U}))$ [10, Thm 3.7, p. 444]. An analogous bound is even easier to verify for the $\mathsf{W}_{1}$ result (3.1). Thus, the $\mathcal{G}^{\star}$-dependent solution of the proposed problem (4.7) over $\mathsf{P}$ (and hence also the solution of Eqn. (4.1)) is always at least as good as the barycenter, which only depends on $\mathsf{P}$ and $\mathbb{Q}$.
@@
and $f\colon\mathcal{U}\to\mathbb{R}$. The next result differentiates $F$; the proof is in SM D.3.
@@
Taking the gradient of $u\mapsto DF(\nu)(u)$ in (4.9) yields the Wasserstein gradient, which can be utilized in Wasserstein gradient flow discretizations; see Rmk. 4.4. In the case of parametric distributions, combining Prop. 4.9 with the chain rule yields gradient descent schemes similar to those introduced in Sec. 4.1. Again, the derived gradients are exact and require only discretization for their numerical evaluation. For further details, see Cor. D.7 and Lem. D.8 and its special case (Rmk. D.9 in SM D.3).
@@ 5 Numerical results
This section implements Alg. 1 (the bilevel gradient descent algorithm) for function approximation using kernel methods and Alg. 2 (the alternating minimization algorithm) for several operator learning tasks using DeepONets. Comprehensive experiment details and settings are collected in SM F.
@@
We instantiate a discretization of Alg. 1 on several function approximation benchmarks. The test functions $\mathcal{G}^{\star}\in\{g_{i}\}_{i=1}^{4}$ each map $\mathbb{R}^{d}$ to $\mathbb{R}$ and are comprised of the Sobol G ($g_{1}$) and Friedmann functions ($g_{2},g_{3}$), and a kernel expansion ($g_{4}$). We use the kernel framework from Sec. 4.1 with $\mathcal{H}$ the RKHS of the squared exponential kernel with a fixed scalar bandwidth hyperparameter. Also, $\mathbb{Q}$ is an empirical measure with $10$ atoms centered at fixed Gaussian measures realized by standard normally distributed means and Wishart distributed covariance matrices. A cosine annealing scheduler determines the gradient descent step sizes. The training distribution $\nu_{\vartheta}\coloneqq\mathcal{N}(m,\mathcal{C})$ is parametrized by the pair $\vartheta=(m,\mathcal{C})$, which we optimize using Alg. 1.
@@
Table 1: Function approximation with bilevel Alg. 1. Reported values are mean $\mathsf{Err}$ for kernel regressors trained with $1024$ labeled data samples from the optimized $\nu_{\vartheta}$ (ours) from Alg. 1 with $1000$ iterations vs. other adaptive (aCoreSet) and nonadaptive (Normal, Barycenter, Mixture, Uniform, nCoreSet) sampling distributions. Two standard deviations from the mean $\mathsf{Err}$ over $10$ independent runs is also reported. Also see Fig. 10.
$g_{1}~(d=2)$ $g_{2}~(d=5)$ $g_{2}~(d=8)$ $g_{3}~(d=4)$ $g_{3}~(d=5)$ $g_{4}~(d=10)$
Ours $\textbf{0.040}\pm 0.010$ $\textbf{0.169}\pm 0.007$ $\textbf{0.556}\pm 0.007$ $\textbf{0.276}\pm 0.019$ $558\pm 0.009 0.0$$025\pm 0.000$
Normal $0.808\pm 0.015$ $0.811\pm 0.010$ $0.949\pm 0.003$ $0.937\pm 0.005$ $947\pm 0.003 0.0$$029\pm 0.001$
Barycenter $0.546\pm 0.025$ $0.514\pm 0.012$ $0.782\pm 0.005$ $0.771\pm 0.008$ $784\pm 0.010 0.0$$021\pm 0.001$
Mixture $0.308\pm 0.011$ $0.372\pm 0.015$ $0.712\pm 0.011$ $0.471\pm 0.037$ $658\pm 0.011 0.0$$022\pm 0.001$
Uniform $0.983\pm 0.001$ $0.971\pm 0.002$ $0.993\pm 0.000$ $0.976\pm 0.000$ $1.000\pm 0.000$ $0.121\pm 0.001$
nCoreSet $0.185\pm 0.000$ $0.247\pm 0.002$ $0.622\pm 0.002$ $0.295\pm 0.002$ $\textbf{0.494}\pm 0.002$ $0.098\pm 0.012$
aCoreSet $0.501\pm 0.060$ $0.538\pm 0.028$ $0.782\pm 0.018$ $0.793\pm 0.048$ $848\pm 0.034 0.0$$\textbf{0.013}\pm 0.0$$0$
Tab. 1 shows that models trained on the final iterate $\nu_{\vartheta}$ of Alg. 1 nearly always outperforms those trained on adaptive or nonadaptive baseline distributions in terms of the root relative average OOD squared error $\mathsf{Err}\coloneqq(\mathscr{E}_{\mathbb{Q}}(\widehat{\mathcal{G}}^{(\nu)})/\mathscr{E}_{\mathbb{Q}}(0))^{1/2}$. Nonadaptive (nCoreSet) and adaptive (aCoreSet) pool-based coresets from active learning [71] slightly outperform $\nu_{\vartheta}$ on the $g_{3}$ $(d=5)$ and $g_{4}$ $(d=10)$ test problems, respectively. Here, the fixed pool of size $5000$ consists of the empirical samples of samples from $\mathbb{Q}$ available to Alg. 1. Overall, the performance gap between our bilevel method and the baselines can close depending on the complexity of $\mathbb{Q}$ or $\mathcal{G}^{\star}$. Fig. 2 specializes to the $\mathcal{G}^{\star}=g_{1}$ case. Although fitting the model on data sampled from the precomputed optimal $\nu_{\vartheta}$ uniformly outperforms baseline distribution sampling for large enough $N$, when corrected for the “offline” cost of finding the optimal $\nu_{\vartheta}$, the gains are more modest (Fig. 2, right). Moreover, Fig. 2 (center) shows that the nCoreSet method based on raw kernel feature maps performs better than our $\nu_{\vartheta}$ and empirical $\mathbb{Q}$ mixture for very small sample sizes, but eventually saturates to the mixture error level as $N$ approaches the pool size. In contrast, the error of $\nu_{\vartheta}$ continues to decrease.
@@
Figure 2: Alg. 1 applied to ground truth $g_{1}\colon\mathbb{R}^{2}\to\mathbb{R}$. (Left) Evolution of $\mathsf{Err}$ over $1000$ iterations ($N=250$ samples per step) of gradient descent. (Center) $\mathsf{Err}$ of model trained on $N$ samples from optimized $\nu_{\vartheta}$ (ours) vs. initial normal $\mathcal{N}(m_{0},I_{2})$, empirical $\mathbb{Q}$ $\mathsf{W}_{2}$-barycenter, empirical $\mathbb{Q}$ mixture, $\mathrm{Unif}([0,1]^{2})$, and two pool-based coresets. (Right) Same as center, except incorporating the additional function evaluation cost incurred from Alg. 1. Shading represents two standard deviations away from the mean $\mathsf{Err}$ over $10$ independent runs.
@@
We apply the alternating scheme Alg. 2 to several important operator learning problems from the physical sciences. Let $\Omega\subset\mathbb{R}^{2}$ denote a bounded domain with boundary $\partial\Omega$. Consider the elliptic PDE $\nabla\cdot(a(x)\nabla u(x))=0$, where $x\in\Omega$, $a=a(x)>0$ is the conductivity, and $u=u(x)$ is the PDE solution. Given a mean-zero Neumann boundary condition (BC) $g=a\frac{\partial u}{\partial\textbf{n}}|_{\partial\Omega}$, one can obtain a unique PDE solution $u$ and the resulting Dirichlet data $f\coloneqq u|_{\partial\Omega}$. For a fixed conductivity $a$, we define a Neumann-to-Dirichlet (NtD) map $\Lambda_{a}^{-1}\colon g\mapsto f$, which is also used in the EIT problem [29]; see also Sec. 1. On the other hand, the map $\mathcal{G}\colon a\mapsto u$ with a fixed Neumann or Dirichlet BC (and nonzero source) is the Darcy flow parameter-to-solution operator, which plays a crucial role in geothermal energy extraction and carbon storage applications. Another important operator maps the scattering coefficient to the spatial-domain density in the radiative transport equation, which has key applications in optical tomography and atmospheric science. Last, we consider the viscous Burgers’ equation with small viscosity; it models steep shock-like phenomena in fluids. Learning surrogate operators for these mappings can enable rapid simulations for efficient inference, design, and control.
@@
Figure 3: (Left) Matrix approximation of NtD map. (Center) After $80$ independent runs, decay of the average relative OOD error of the model when trained on the optimal distribution identified at each iteration of Alg. 2; a 9595% confidence interval of the true relative OOD error is provided at each iteration. (Right) Decay of average AMA loss defined in (4.7) vs. iteration relative to the same loss at initialization.
@@
Radiative transport operator learning. We apply a nonparametric particle-based version of AMA to find the optimal training distribution for learning the solution operator of the radiative transport equation. We vary the Knudsen number which corresponds to different physical regimes. Fig. 5 shows how the model trained on the optimal empirical measure compares to the initial model and a benchmark model trained on the empirical $\mathbb{Q}$ mixture. When the sample size is $N=120$ particles, particle-based AMA reduces the average relative OOD error by 88% compared to its initial value for a Knudsen number of 8.
(a) $\varepsilon=1/8$
(b) $\varepsilon=2$
(c) $\varepsilon=8$
Figure 5: Relative OOD error vs. sample size NN for learning the radiative transport solution operator. Results are shown for Knudsen numbers $\varepsilon\in\{1/8,2,8\}$. For each $N$, each panel displays 95% confidence intervals over 10 trials for three DeepONet models trained for 5000 epochs: the initial model, the model after particle-based AMA, and a benchmark trained the test distribution mixture $\nu_{\mathbb{Q}}\coloneqq\frac{1}{3}\sum_{k=1}^{3}\nu_{k}^{\prime}$.
Burgers’ equation operator learning. We apply AMA to learn the operator that maps an initial condition $u({\,\cdot\,},0)$ to the corresponding solution $u({\,\cdot\,},1)$ of the viscous Burgers’ equation. The left and center panels of Fig. 6 illustrate an example input-output pair for this operator. We compare AMA against two pool-based active learning methods: Query-by-Committee (QbC) and CoreSet [71]. Using a fixed pool, we first compute the optimal training distribution from AMA and then draw $N$ samples from it to train the DeepONet model. The same pool is used to obtain models trained with each active learning strategy. The right panel of Fig. 6 reports the relative OOD performance as $N$ increases. The AMA-trained model performs comparably to the active learning baselines and, in the mid-range of sample sizes, exceeds their performance. When the sample size is small, all methods begin with randomly selected points, leading to similar performance. For large sample sizes, the training sets seem to overlap, causing the methods to converge. In the intermediate regime, however, our model consistently outperforms the other two approaches. Moreover, AMA provides the additional advantage of generating arbitrarily many high-quality training samples beyond the fixed pool, yielding not only a stronger learned operator but also a robust training distribution.
Figure 6: (Left) Example of a random initial condition $u({\,\cdot\,},0)$ for the Burgers’ equation. (Center) Corresponding solution $u({\,\cdot\,},1)$ of the Burgers’ equation at time 11 for the given initial condition. (Right) Relative OOD error vs. sample size NN for learning the mapping from the initial condition $u({\,\cdot\,},0)$ to the solution $u({\,\cdot\,},1)$. After computing the optimal training distribution using a fixed data pool, the model labeled “AMA” is trained on $N$ samples drawn from this distribution. This model is compared with two models trained using pool-based active learning methods, QbC and CoreSet. All models share the same DeepONet architecture and differ only in the training data received. The experiment is repeated 10 times, and the plot shows the 95% confidence interval for the expected relative OOD error.